Jack invests $1000 at a certain annual interest rate, and he invests another $2000 at an annual rate that is one-half percent higher. If he receives a total of $190 interest in 1 year, at what rate is the $1000 invested?
Answer by Boreal(15235)

interest rate 1=x
interest rate 2=0.005+x
(1000*x)+(2000)*(x+0.005)=190
1000x+2000x+10=190
3000x=180
x=0.06 or 6%
x+0.005=0.065 or 6.5%
1000*6%=$60 per year ANSWER IS 6%
2000*6.5%=$130per year
$190 per year.